OpenStudy (anonymous):

How do I use log here? 75000 * .96^x = 50000 * 1.07^x x = ?

OpenStudy (saifoo.khan):

\[0.96^x = \frac{50000}{75000} * 1.07^x\] \[0.96^x = \frac23 * 1.07^x\]

OpenStudy (saifoo.khan):

Radar will continue from here.

OpenStudy (radar):

Hey I was in the learning mode, keep on please.

OpenStudy (saifoo.khan):

Lol, i forgot how to solve these. Hold a sec. let me get back my memory.

OpenStudy (radar):

I don't think I ever knew lol

OpenStudy (saifoo.khan):

we have to "suppose a particular value for this) Let me get my register real fast.

OpenStudy (saifoo.khan):

Called 911 for help.

OpenStudy (bahrom7893):

take ln on both sides

OpenStudy (saifoo.khan):

Still stuck..

OpenStudy (bahrom7893):

75000 * .96^x = 50000 * 1.07^x ln(75000 * .96^x) = ln(50000 * 1.07^x)

OpenStudy (bahrom7893):

Saifoo please let me finish

OpenStudy (bahrom7893):

ln75000+ln0.96^x = ln50000+ln1.07^x

OpenStudy (bahrom7893):

xln0.96 - xln1.07 =ln50000-ln75000

OpenStudy (bahrom7893):

x(ln0.96-ln1.07) = ln50000-ln75000

OpenStudy (bahrom7893):

x = (ln50000-ln75000)/(ln0.96-ln1.07) finish the algebra

OpenStudy (anonymous):

why the natural log, could you use regular log?

OpenStudy (bahrom7893):

lol u can i guess, it's just that im so used to natural logs.

OpenStudy (bahrom7893):

In calc u barely use any other logs. I'm really used to natural logs, lol it feels natural.

OpenStudy (anonymous):

so if I substituted log in for each ln it would give me the same answer?

OpenStudy (bahrom7893):

it should

OpenStudy (anonymous):

ahah that's really punny stuff

OpenStudy (radar):

I will attempt and then saifoo.khan can correct me. Continuing with the last equation. Dividing both sides by 0.96^x getting 1=2/3 * 1.07^x/.96^x multiplying both sides by 3/2, we can now write it as 3/2=(1.07/.96)^x simplfying 1.5=1.1146^x log 1.5= x log 1.1146 .1761 = x *.04712 x=3.7

OpenStudy (anonymous):

what if the question had said 10000 + .96^x = 1.07^x

OpenStudy (bahrom7893):

hmm then throw 0.96^x to the right and take log on both sides.

OpenStudy (bahrom7893):

wait a second..

OpenStudy (radar):

It is probably closer to 4 as I rounded off the log values.

OpenStudy (radar):

Wolfram shows 4.

OpenStudy (bahrom7893):

btw idk what happens if it's pluses, logs wont work, ull need a calculator.

OpenStudy (anonymous):

oh I see what i did wrong, thanks for showing me how guys

OpenStudy (bahrom7893):

http://www.polyvore.com/cgi/img-thing?.out=jpg&size=l&tid=31489685 lol sorry, but ur last post asked for this hahahaaha

OpenStudy (anonymous):

rofl